### v2

# Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection

Abstract
The spread of fake news negatively impacts individuals and is regarded as a significant social challenge that needs to be addressed. A number of algorithmic and insightful features have been identified for detecting fake news. However, with the recent LLMs and their advanced generation capabilities, many of the detectable features (e.g., style-conversion attacks) can be altered, making it more challenging to distinguish from real news. This study proposes adversarial style augmentation, AdStyle, to train a fake news detector that remains robust against various style-conversion attacks. Our model’s key mechanism is the careful use of LLMs to automatically generate a diverse yet coherent range of style-conversion attack prompts. This improves the generation of prompts that are particularly difficult for the detector to handle. Experiments show that our augmentation strategy improves robustness and detection performance when tested on fake news benchmark datasets.
Adversarial Style Augmentation via Large Language Modelfor Robust Fake News Detection
Sungwon Park1, Sungwon Han1, Meeyoung Cha2 1School of Computing, KAIST, Daejeon, Republic of Korea 2MPI for Security and Privacy, Bochum, Germany {psw0416, lion4151}@kaist.ac.kr, mia.cha@mpi-sp.org
1 Introduction
With the widespread use of the internet and the emergence of various social media platforms, people have begun to freely share information they know and their own creative stories. However, the ease of sharing and consuming information beyond traditional news organizations has also significantly contributed to the spread of fake news Cha et al. (2021); Fisher et al. (2016); Treen et al. (2020). Fake news is created to provide readers with false information as propaganda or to guide public perception in a desired direction for intentional objectives Lazer et al. (2018); Yi et al. (2024). The spread of such false information has had profound negative impacts on individuals and society, becoming a major social challenge that needs to be addressed.
Manually determining the authenticity of all information on the internet is extremely costly in terms of time and resources. Thus, prior advances have mainly focused on developing automated fake news detectors using machine learning techniques Pérez-Rosas et al. (2018); Reis et al. (2019); Shu et al. (2019). For example, several studies have extracted sentiment-related or political features commonly found in fake news, either through human-crafted Potthast et al. (2018); Rashkin et al. (2017) or data-driven methods Reis et al. (2019); Shu et al. (2019), and used them for detection. With the advent of language models (e.g., GPT, BERT) Kenton and Toutanova (2019); Radford et al. (2019), efforts have been made to learn the textual style of fake news by extracting sentence embeddings to train detectors Chen et al. (2021); Yang et al. (2022).
While these extracted textual styles or human-crafted features have proven effective in identifying fake news, they allow attackers to bypass detection. In particular, the advent of large language models (LLMs) Chowdhery et al. (2022); OpenAI (2023) has made it possible to easily and automatically paraphrase sentences in a user-desired direction through prompts (e.g., "change the given text to an objective and professional style"), making it more difficult to distinguish between AI-generated fake news and real news (i.e., style-conversion attacks) Koenders et al. (2021); Zhou et al. (2019). Recent literature has proposed strategies to augment the styles of input text via some manually defined style-conversion prompts, which are taken by LLM for paraphrasing, but it remains challenging to address all types of prompts that attackers might use to deceive detectors Wu and Hooi (2023).
In this study, we propose adversarial style augmentation, AdStyle, to train a robust fake news detector that can withstand various style-conversion attacks by attackers. In contrast to earlier work, which used predefined style-conversion prompts as detector-agnostic augmentations, AdStyle tries to find prompts for adversarial style augmentations that are specific to the detector, which makes predictions uncertain. This means that our augmentations add noise to the style features in the direction of the detector’s decision boundary, while maintaining the text’s content integrity. Specifically, to search for and optimize prompts for LLM that are adversarial to the detector, we introduce an automated prompt engineering technique Yang et al. (2023). By providing LLMs with style-conversion prompts and the detector’s performance under these augmentations, LLM can infer patterns between the prompts and performance, enabling the search for the most adversarial prompts.
AdStyle proceeds as follows: To generate augmented samples, we first select a random subset of the dataset and apply a pool of style-conversion prompts to it. Then, each augmentation set created by different prompts is fed into the detector, and the score based on the AUC between the predictions and the ground-truth labels is measured. The prompt-score pairs are provided to the LLM, which uses this information to generate new style-conversion prompt candidates. From these candidates, we select the top-$k$ prompts that are diverse and make the detector’s predictions most uncertain without significantly altering the original content. These selected prompts are applied to the entire dataset and used for detector training. The chosen prompts are added to the prompt pool, which is repeated over several training rounds.
We experimented with existing fake news benchmark datasets under various style-conversion attack scenarios. As a result, our augmentation strategy demonstrated higher robustness and detection performance compared to previous methodologies. Additionally, our approach is scalable to different attack strategies by adding new attack prompts to the style-conversion prompt pool. We plan to release our code after publication.
2 Related Work
2.1 Automated Detection of Fake News
Manually detecting fake news among the vast amount of content on the internet is intractable. Consequently, many studies have focused on leveraging machine learning techniques to automatically detect fake news Pérez-Rosas et al. (2018); Reis et al. (2019); Shu et al. (2019). For example, supervised learning methodologies extract textual features from fake news texts using benchmark datasets and ground-truth labels Reis et al. (2019). These textual features can include deep features based on artificial neural networks Reis et al. (2019); Shu et al. (2019) as well as manually defined features such as sentiment and political bias Potthast et al. (2018); Rashkin et al. (2017). Other approaches include domain adaptation methods to enhance detection generalizability across various domains and topics Mosallanezhad et al. (2022); Nan et al. (2022), and knowledge-based methods that rely on external data to distinguish false information Dun et al. (2021). With the advent of LLMs, new methods have emerged that utilize LLMs’ prior knowledge to identify fake news Hu et al. (2024), including research on detecting machine-generated fake news Chen and Shu (2023).
In this work, we aim to prevent malicious attackers from deceiving automated fake news detectors through text modifications such as paraphrasing. The proposed method is an augmentation strategy to enhance robustness against text style perturbations as an add-on to existing detection models. Our contribution is agnostic to the design of the detection model.
Figure 1: Illustration of how AdStyle works. Our model consists of multiple rounds of training. In each round, the model uses the style-conversion prompt and prediction confusion score from the previous round to generate prompt candidates that can maximize the confusion of the detector model. Subsequently, considering aspects such as adversarialness, coherency, and diversity, a subset of the training dataset is used to select the top-$k$ prompts, which are finally used as augmentations in training.
2.2 Attack on Fake News Detection
To test the robustness of fake news detection, various attack methods have been studied Koenders et al. (2021); Zhou et al. (2019). These include injecting misinformation by changing the order of subjects and objects or causes and effects while maintaining the textual features used for detection Koenders et al. (2021). Other methods involve creating fact distortions by altering or exaggerating words related to people, time, or places while preserving the sentence structure Zhou et al. (2019). Additionally, approaches that use the text generation capabilities of LLMs to change the style of sentences have been proposed Wu and Hooi (2023).
Our method aims to perform adversarial training of the detection model by generating augmentations that perturb the text’s style while preserving the content as much as possible. This approach reduces the impact of spurious style features on the detection model, making it more robust against such attacks.
2.3 Prompt Engineering for LLM
Training with extensive text corpora, LLMs have demonstrated their utility across various domain tasks Chowdhery et al. (2022); OpenAI (2023); Pang et al. (2024). To better harness the prior knowledge and reasoning abilities of LLMs, strategies for providing appropriate input prompts have also been researched. For instance, in-context learning methods involve providing examples of the desired input-output format to guide the LLM in producing the correct output Dong et al. (2022). Additionally, engineering techniques such as chain-of-thought prompting, which includes additional reasoning steps in the responses, have been developed to enhance the reasoning abilities of LLMs Wei et al. (2022). Recently, methods for automatically finding the most suitable prompts for a given task, known as automated prompt engineering, have shown promising results Yang et al. (2023); Zhou et al. (2022).
In this paper, we adopt automated prompt engineering techniques to identify adversarial prompts that most effectively confuse the fake news detector. This process replaces the traditional gradient descent method used in the image domain to find adversarial noise, enabling the creation of adversarially augmented versions of input text via LLM.
3 Method
3.1 Overview
Let $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$ be a dataset containing news $\mathbf{d}_{i}$ and the corresponding ground-truth binary veracity label $y_{i}$ (indicating whether the news is true or fake). Each news item $\mathbf{d}_{i}$ is composed of natural language-based text. This study aims to train a language model based fake news detector $f$ using the labeled dataset to predict the veracity labels. Our main goal is to develop a detector that remains robust even when an attacker perturbs the textual style, such as the order and format, while preserving the meaning of the sentences.
Figure 1 illustrates how our model works. AdStyle generates style-conversion prompts and performs augmentation over multiple rounds. Each round includes a following process. Firstly, leveraging the reasoning ability of LLM, adversarial style-conversion prompt candidates are generated. These style-conversion prompts contain instructions on how to transform the given sentences and are used as inputs to the LLM along with the original sentences for conversion. Inspired by automated prompt engineering Yang et al. (2023), the style-conversion prompts and detector prediction score pairs used in previous rounds are included as in-context demonstrations to guide the LLM in searching prompt candidates that maximally confuse the detector (Section 3.2). Next, a subset of the dataset is selected, and these discovered candidates are applied to perform conversions. The converted samples are then evaluated to determine how much they confuse the detector. From these candidates, the top-$k$ prompts that are diverse and maintain the original content’s meaning while most effectively confusing the detector are selected (Section 3.3). These selected prompts are used as an augmentation method to train the detection model in the current round. We describe each step’s details below.
3.2 Generating Adversarial Style-Conversion Prompts with LLM
The style-conversion prompts we aim to generate are instructions that perturb only the textual style, such as the structure or format of the sentence, while preserving the content of the given input text. For example, an instruction like “Rewrite the following article in an objective and professional tone” can be used to transform the stylistic features of fake news to resemble real news. However, if such instructions are heuristically defined, they may not align with the actual decision boundary of the detector, reducing the training efficiency. Furthermore, if the instructions remain fixed throughout the training, the detector may memorize these conversion patterns, resulting in overfitting.
Therefore, instead of predefining and fixing the instructions for conversion, we aim to find adversarial conversion prompts that add noise in the direction of the decision boundary of current state of detector, maximizing the confusion in the detector’s predictions. This approach is similar to adversarial training commonly used in the computer vision domain Madry et al. (2018), which enhances robustness against slight perturbations in the input. However, for textual data, the discrete nature of the input makes it challenging to add noise directly through the detector’s gradient as in the image domain. To address this, we measure the confusion in prediction by conversion prompts (i.e., prediction confusion score), and feed this prompt-score pair to LLM for generating adversarial prompts, motivated from automated prompt engineering techniques.
Now you will help me minimize a fake news detector performance value with style transfer attack instruction. I have some style transfer attack instructions and the fake news detector performance value using those instructions. The pairs are arranged in descending order based on their function values, where lower values are better. Input: [Style-Conversion Prompt Example] Score: [Prediction Confusion Score] Input: … … Give me a new style transfer attack instruction that is different from all pairs above, and has a performance value lower than any of the above.
Figure 2: Prompt for generating adversarial style-conversion prompts. The text in blue represents the score trajectory component, while the remaining text represents the problem description component.
Figure 2 shows an example of the LLM input used to generate adversarial style-conversion prompts. The LLM input consists of the problem description and the score trajectory components.
Problem description component.
This component includes the problem description, the objective, and constraints on the response necessary for generating style-conversion prompts. For example, a sentence like “Minimize a fake news detector performance value with style transfer attack instruction” informs the LLM of the intent behind the conversion prompt.
Score trajectory component.
Previous work has shown that LLMs can learn patterns from in-context demonstrations provided as input Dong et al. (2022); Yang et al. (2023). This component leverages this ability by providing previous round’s style-conversion prompts and their corresponding prediction confusion scores in the form of in-context demonstrations [^1] In the first round, a predefined set of prompts is used.. The score is measured by selecting a subset $\mathcal{B}$ from the entire training dataset $\mathcal{D}=\{(\mathbf{d}_{i},y_{i})\}_{i=1}^{N}$, applying a conversion prompt $c$ to create a new set $\mathcal{B}^{c}=\{(\mathbf{d}^{c}_{i},y_{i})\}_{i=1}^{M}$, where $\mathbf{d}^{c}_{i}=\text{Convert}(c,\mathbf{d}_{i}),N>>M$, and then measuring the AUC score between the predictions and the ground-truth labels when $\mathcal{B}^{c}$ is fed into the detector. Specifically, the score of a conversion prompt $s_{c}$ is defined as:
$s_{c}=|0.5-\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})|.$ (1)
A lower score indicates a higher level of prediction confusion, implying that the conversion prompt has caused the detector’s predictions to become random with respect to the labels. Finding conversion prompts that cause high confusion (i.e., low score) suggests that the detector has not yet learned to handle those stylistic features, and the conversion prompts have placed the samples near the detector’s decision boundary, making them difficult to distinguish. By showing the LLM these in-context demonstrations, it can generate conversion prompts that differ from past ones while maximizing confusion (i.e., minimizing the score). Note that we avoid selecting conversion prompts that flip the detector’s original predictions (i.e., AUC <<much-less-than$<$ < 0.5), as these often disrupt the content along with the stylistic features, which is undesirable. We extract $S$ style-conversion prompts at a time using the input described above.
3.3 Selecting Top-$k$ Adversarial Prompts
Using all the style-conversion prompt candidates generated by the LLM can be computationally intensive, and not all prompts may be suitable for augmentation. For example, the set of style-conversion prompts used for augmentation should each provide adversarial perturbations that confuse the detector (i.e., adversarialness). Additionally, while altering the textual style, the prompts should not change the meaning of the sentences to prevent label noise (i.e., coherency). Furthermore, the more diverse the set of conversion directions covered by the prompts, the more efficient the augmentation process (i.e., diversity).
To select a set of style-conversion prompts that satisfies these three criteria, we propose a selection strategy. Given a conversion prompt $c$, we first extract embedding vectors for the input texts in both the training subset $\mathcal{B}$ and the subset $\mathcal{B}^{c}$ converted by $c$ using a large language model $g$ such as BERT. We then compute the average embedding vectors for each subset and calculate the vector difference $\mathbf{z}_{c}$.
$\mathbf{z}$ $=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{d}_{i}\in\mathcal{B}}{g(\mathbf{d}_{i})},\ \ \ \mathbf{z}\textquoteright=\frac{1}{|\mathcal{B}\textquoteright|}\sum_{\mathbf{d}^{c}_{i}\in\mathcal{B}\textquoteright}{g(\mathbf{d}^{c}_{i})}$
$\mathbf{z}_{c}$ $=\mathbf{z}\textquoteright-\mathbf{z}$ (2)
Here, $\mathbf{z}_{c}$ represents the average change in embedding direction due to the conversion prompt $c$. We then calculate the adversarialness scale $s^{c}_{\text{adv}}$ and coherency scale $s^{c}_{\text{coh}}$ for each conversion prompt $c$, and rescale $\mathbf{z}_{c}$ accordingly (i.e., $\mathbf{\hat{z}}_{c}=\mathbf{z}_{c}\times s^{c}_{\text{adv}}\times s^{c}_{\text{coh}}$). Finally, we use the $k$-means++ initialization method Arthur et al. (2007) on these rescaled vectors to select $k$ prompts. The $k$-means++ initialization method helps select a diverse set of prompts that are adversarial and coherent, by choosing samples that are as far apart as possible Ash et al. (2019). The details of each scale are described below.
Adversarialness scale.
To measure how adversarial a given style-conversion prompt is to the detector, we define a new adversarialness scale similar to the confusion score defined in the previous section. Given the converted batch $\mathcal{B}^{c}$ by conversion prompt $c$, the adversarialness scale $s^{c}_{\text{adv}}$ is defined as $-1.8\cdot|\text{AUC}(\{y_{i}\}_{i=1}^{M},\{f(\mathbf{d}^{c}_{i})\}_{i=1}^{M})-0.5|+1$. This value increases as the AUC approaches 0.5, indicating that the prediction is more random. The coefficient 1.8 ensures that the scale ranges between 0.1 and 1, preventing it from being zero.
Coherency scale.
To verify that the converted text retains the same content as the original text, we check the similarity in meaning between the text pairs using an LLM. Specifically, we create sample pairs from $\mathcal{B}$ and the converted subset $\mathcal{B}^{c}$, and inquire the LLM about the percentage of pairs that it considers to have the same meaning. This percentage is used as the coherency scale $s^{c}_{\text{coh}}$. Like the adversarialness scale, this value is rescaled to range between 0.1 and 1.
Finally, the selected style-conversion prompts are applied to the input texts of the entire dataset $\mathcal{D}$ to create augmented samples. These augmented samples are then used alongside the original samples to train the detector $f$. The detector is trained using binary cross-entropy loss. These prompt generation and selection processes are repeated over multiple training rounds.
4 Experiment
Table 1: Performance comparison with AdStyle in two different scenarios—Attack, where style-conversion attacks are performed, and Clean, where no attack is performed—across three fake news datasets. For the attack scenario, we report the average AUC of four style-conversion attacks. The best results are marked in bold. Our model demonstrates a significant performance improvement over all text-based fake news detectors in both style-conversion attack and clean scenarios.
Attack Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6114 0.6904 0.8548 0.6920 0.7876 0.8453 0.8185 0.8674 0.8741
UDA 0.6241 0.7696 0.8564 0.7381 0.7865 0.8591 0.8615 0.9028 0.9297
RADAR 0.7218 0.7399 0.8571 0.7583 0.8170 0.8616 0.8535 0.8047 0.9086
ENDEF 0.6376 0.7579 0.8134 0.7405 0.7870 0.8615 0.8234 0.8950 0.8835
SheepDog 0.6525 0.8234 0.9009 0.7498 0.8357 0.8669 0.8926 0.9188 0.9630
AdStyle 0.7833 0.8919 0.9399 0.8134 0.8389 0.8721 0.9224 0.9531 0.9716
Clean Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.7393 0.8397 0.9355 0.7096 0.8104 0.8645 0.9311 0.9682 0.9892
UDA 0.7404 0.8783 0.9422 0.7422 0.8022 0.8666 0.9365 0.9724 0.9899
RADAR 0.7607 0.8495 0.9314 0.7593 0.8170 0.8630 0.9446 0.9773 0.9817
ENDEF 0.7776 0.8823 0.9294 0.7592 0.7991 0.8738 0.9234 0.9556 0.9871
SheepDog 0.7248 0.8229 0.9394 0.7490 0.8411 0.8641 0.9144 0.9459 0.9785
AdStyle 0.8996 0.9280 0.9460 0.8251 0.8493 0.8797 0.9509 0.9849 0.9889
Table 2: Performance comparison of ablations on the Politifact dataset. The results show the impact of style-conversion attacks using four different publishers (i.e., CNN, The New York Times, The Sun, and National Enquirer) and a clean scenario (i.e., Clean) where no attack is performed. Any modification or removal of model components leads to decreased performance.
Model CNN The New York Times The Sun National Enquirer Clean
Vanilla 0.8127 0.8687 0.8789 0.8591 0.9355
Random Selection 0.9075 0.9409 0.9355 0.9365 0.9412
Adv. only Selection 0.9035 0.9343 0.9318 0.9288 0.9405
w/o Uncertainty 0.9039 0.9333 0.9320 0.9402 0.9473
w/o Similarity 0.9193 0.9315 0.9297 0.9430 0.9409
w/o Score trajectory 0.8199 0.8917 0.9124 0.9066 0.9372
Full Components 0.9174 0.9444 0.9520 0.9460 0.9460
We evaluate the robustness of AdStyle under diverse style-conversion attacks across multiple datasets, comparing it with contemporary baselines. Then, we analyze the impact of model components on overall performance. Lastly, we conducted a qualitative analysis to investigate the characteristics of the style-conversion prompts generated by the LLM. Due to limited space, additional analysis, such as the performance evaluation over wider variety of paraphrasing attacks and comparisons with LLM-based zero-shot and in-context learning baselines, has been included in the Appendix.
4.1 Performance Evaluation
Dataset.
Our experiments use three real-world fake news benchmark datasets. We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al. (2020), which focus on political claims and celebrity rumors, respectively. Additionally, we incorporate Constraint Felber (2021), a dataset specifically addressing COVID-19 related social media posts. Each dataset is randomly split into an 80% training set and a 20% test set. Detailed dataset statistics are provided in Table 3 of the Appendix.
Attack settings.
To assess robustness against style conversion attacks, we employ LLM-empowered techniques to reframe the test set using a variety of style conversion prompts, as illustrated in Figure 3. Following the original literature Wu and Hooi (2023), we use four well-known daily news sources as [publisher name]: CNN, The New York Times, The Sun, and National Enquirer. CNN and The New York Times are recognized for their reputable journalism, while The Sun and National Enquirer are characterized by their tabloid style. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences, with the temperature set to 0 and the top-p value set to 1 by default. We conduct experiments using 10%, 25%, and 100% of the complete dataset to observe the effect of augmentation across different dataset sizes for training. AUC is used as our evaluation metric.
Rewrite the following article using the style of [publisher name]: [news article]
Figure 3: Prompt for style conversion. The “publisher name" part will be filled with the name of a representative publisher (e.g., newspaper or journal), and the “news article" part will contain the original news text.
Baselines.
We have implemented several existing text-based fake news detection strategies as baselines: (1) Vanilla, a text-based fake news detector using conventional binary cross entropy objective; (2) UDA, introduces consistency regularization objective between original text and diverse augmented variations. Xie et al. (2020); (3) RADAR, utilizes an adversarially trained paraphraser to generate augmented version of input sentences Hu et al. (2023); (4) ENDEF, mitigates entity bias in fake news data through causal learning Zhu et al. (2022); (5) SheepDog, introduces predefined style-conversion prompts to augment the styles of input text via LLM Wu and Hooi (2023). We follow the original paper’s setting and details for baseline implementations.
Implementation details.
All models are evaluated under uniform experimental conditions to ensure fair comparison. This consistency extends to the choice of backbone network, optimizer, and learning rate. We utilize OpenAI’s GPT-3.5-Turbo model for reframing input sentences and measuring coherency, with the temperature set to 0 and the top-p value set to 1 by default. Our training process for the detector is conducted over 10 rounds, with one training epoch per round. When evaluating the style-conversion prompts, we randomly selected 30 samples from the training dataset to apply augmentation (i.e., $M=30$). In each round, 30 prompt candidates were generated by the LLM (i.e., $S=30$), from which 3 were chosen for augmentation (i.e., $k=3$) by our selection strategy. The training utilizes the AdamW optimizer with a learning rate of 1e-5 and a batch size of 8. For measuring diversity, the text embeddings are generated using a pre-trained BERT-based uncased model from HuggingFace Transformers. Two V100 GPUs were utilized for all experiments.
Result.
Table 1 compares the performance of detector algorithms in both clean scenarios, where no attack is performed, and adversarial scenarios, where style-conversion attacks are applied. Due to space limitations, the average AUC results under the four different style-conversion attacks are reported. Detailed results for each prompt can be found in the Appendix (See Table 6 in the Appendix). According to the results, AdStyle consistently outperforms the baselines across all cases, including both clean and adversarial scenarios. This demonstrates that our augmentation strategy enhances both the robustness and generalizability of the detector. Notably, our model’s effectiveness is more pronounced when compared to other baselines, especially with smaller datasets. Figure 4 shows the AUC for each style-conversion attack over different rounds. The performance gradually improves as the rounds progress, indicating that continually discovering adversarial augmentations is beneficial.
Figure 4: Performance changes across rounds on the PolitiFact dataset for four style-conversion attacks. The x-axis represents the training rounds, and the y-axis represents the detector’s AUC. For all attacks, the detector’s performance improved as the rounds progressed.
4.2 In-Depth Performance Analysis
We have demonstrated AdStyle’s effectiveness. In this section, we further examine the contribution of each component and provide a qualitative analysis on our adversarial style-conversion prompts.
Ablation study.
The proposed method integrates two main modules: adversarial style-conversion prompts generation and selection. To evaluate their individual contributions, we conduct experiments where we either remove each component or substituted it with an alternative within the full model. This results in six distinct configurations for analysis: (1) Full Components: Our complete method with all components; (2) Random Selection: The method that randomly select prompts from candidates instead of using our selection strategy; (3) Adversarial only Selection: The method that selects top-$k$ adversarial prompts, not considering diversity and coherence; (4) w/o Adversarialness: The method that omits the adversarialness score in our selection strategy. (5) w/o Coherency: The method that omits the similarity score in our selection strategy. (6) w/o Score trajectory: The method without score trajectory component for the style-conversion prompt generation.
Table 2 demonstrates that omitting any component leads to a decrease in performance for certain style conversions attack. While omitted due to space limitation, it can be confirmed that our sampling strategy converges faster than other sampling strategies (See Figure 8 in Appendix).
(a) Diversity
(b) Adversarialness
(c) Coherency
Figure 5: Qualitative analyses of AdStyle. (a) Histogram of diversity in embedding ($\mathbf{z}_{c}$) of conversion prompts sampled by our method and adversarial only selection. (b) Histogram of adversarialness scores ($s_{adv}^{c}$) for selected and unselected prompts (c) Histogram of coherency scores for selected and unselected prompts
Qualitative analyses.
First, we verified that our sampling strategy effectively selects diverse style-conversion prompts with high adversarialness and coherency. Figure 5(a) visualizes the diversity of prompts selected by our method compared to those chosen solely based on adversarialness scores (i.e., Adversarial-only Selection, the third model in our ablation study). We measure diversity using the average cosine similarity of every pair of selected prompts’ embeddings, $\mathbf{z}_{c}$ (Eq. 2):
$\text{Diveristy}(\mathcal{C})$ $=1-{1\over|\mathcal{C}|}\sum_{(c_{i},c_{j})\in\mathcal{C}}{\text{sim}(\mathbf{z}_{c_{i}},\mathbf{z}_{c_{j}})},$ (3)
where $\mathcal{C}$ is the set of prompt pairs, $\text{sim}(\cdot)$ represents the cosine similarity. The result in the Figure 5(a) indicate that our strategy results in a more diverse set of augmentations compared to selection based on adversarialness alone.
Figure 5(b) and 5(c) illustrate the Adversarialness and Coherency, respectively, of our selected prompts compared to remaining unselected prompts. Adversarialness was measured using the $s_{adv}^{c}$ score (Section 3.3), and coherency was calculated as the cosine similarity between the original text and its augmented version using semantic BERT embeddings Chanchani and Huang (2023). We can also observe that, for both metrics, prompts selected through AdStyle exhibit higher values compared to unselected prompts. This suggests that our strategy effectively selects prompts by considering both adversarialness and coherency.
Next, we qualitatively analyzed the characteristics of our style-conversion prompts. Figure 6 shows examples of prompts generated and selected by AdStyle. Previous work relied on prompts that were more conventional, such as “neutral" or “sensational." However, our prompts included creative phrases beyond typical human suggestions, which proved to be more adversarial to the detector. Currently, we use a fixed format for the initial set of prompts, which results in the generation of prompts with a similar format. We expect that using a more diverse initial set of prompts will allow the LLM to optimize prompts within a wider search space.
Round 0 P1: Rewrite the following article in a nonsensical and absurdly exaggerated tone with a hint of horror P2: Rewrite the following article in a sarcastic and mocking tone P3: Rewrite the following article in a chaotic and disorganized tone Round 1 P1: Rewrite the following article in a haunting and macabre tone with a sense of impending horror and madness P2: Rewrite the following article in a cryptic and enigmatic tone P3: Rewrite the following article in a malevolent and apocalyptic tone with a sense of impending doom and destruction, while also incorporating elements of surrealism and existential dread
Figure 6: Examples of our adversarial style-conversion prompts selected for the PolitiFact dataset
5 Conclusion
This paper presents a robust fake news detection method that withstands various style-conversion and paraphrasing attacks through adversarial style-conversion. Unlike traditional detectors that use predefined, agnostic augmentations, AdStyle employs tailored augmentations that shift samples in the direction of the detector’s current decision boundary using style-conversion prompts, functioning similarly to adversarial noise. Among the various prompt candidates generated by the LLM, we selected an efficient set of prompts for training by considering diversity, coherency, and adversarialness. As a result, we were able to train a detector that exhibits high robustness and generalizability against a wide range of attacks. We believe that our work helps filter fake news and contribute to a better exchange of information in online.
Appendix A APPENDIX
A.1 Data & Implementation Details
Dataset overview.
This paper utilizes a diverse range of datasets sourced from publicly available fake news detection resources, selected for their unique topical focuses. We utilize PolitiFact and Gossipcop, drawn from the FakeNewsNet benchmark Shu et al. (2020), addressing political claims and celebrity rumors respectively. Furthermore, we include Constraint Felber (2021), a dataset focused on COVID-19 related social media posts. Each dataset is randomly divided into an 80% training set and a 20% test set. Detailed statistics for these datasets can be found in Table 3.
Table 3: Statistics of fake news datasets.
Dataset PolitiFact GossipCop Constraint
# of News Articles 774 7,916 8,418
# of Real News 399 3,958 4,406
# of Fake News 375 3,958 4,012
Implementation details.
For consistency, experiments replicating existing baselines maintained fixed settings for learning, backbone network, and other relevant parameters. Augmented variations for UDA were generated using back-translation through German. RADAR employed adversarial training to learn paraphrases, utilizing the maximization of the binary cross-entropy loss of a fake news detector as a reward signal, with the T5-large model serving as the paraphraser. Named entities for ENDEF were extracted using the bert-base-NER model from Hugging Face. For SheepDog, augmentations were generated following the original paper’s methodology, using the prompt format illustrated in Figure 7 and four tones: "objective and professional," "neutral," "emotionally triggering," and "sensational". We also utilized the same four tones and prompt format as our predefined initial prompt set. Two V100 GPUs were utilized for all experiments.
Rewrite the following article in a/an [tone]: [news article]
Figure 7: Prompt for style-conversion. The “tone" part will be filled with the desired tone, and the “news article" part will contain the original news text.
A.2 Further Results on Evaluation.
Result on selection strategies across rounds.
Figure 8 shows the AUC for each selection strategy over different training rounds. Random Selection leads to slow convergence and unstable training, while Adversarial-only Selection, limited by its narrow focus on adversarial prompts, can negatively impact performance in later stages of training due to a lack of diversity in the generated prompts. Notably, without score trajectories, the model struggles to generate prompts that contribute effectively to the training process.
Figure 8: Performance of selection strategies over training rounds on the PolitiFact dataset. The x-axis represents the training rounds, while the y-axis represents the detector’s AUC.
Table 4: Performance comparison of different attack scenarios on the PolitiFact Dataset. (A: Adversarial prompt, B: Summarization prompt, C: In-Context prompt D: Adversarial Paraphraser)
Model A B C D
Vanilla 0.8881 0.8522 0.8896 0.8305
UDA 0.8624 0.8363 0.8924 0.8682
RADAR 0.9096 0.8754 0.9234 0.9007
ENDEF 0.8628 0.7978 0.9009 0.8682
SheepDog 0.9297 0.9205 0.9276 0.8995
Ours 0.9456 0.9416 0.9425 0.9212
Comparison on other possible attack scenarios.
We have conducted additional comparison experiments with more diverse attack scenarios: (1) Adversarial prompt: Given a news article and its label, the prompt instructs the LLM to rewrite the article to evade detection as the given label. (2) Summarization prompt: A prompt instructing the LLM to summarize the news article without incorporating stylistic elements. (3) In-Context prompt: A prompt providing an example of a recent, real CNN article, instructing the LLM to rewrite the given article in the same style. (4) Adversarial Paraphraser: The attack utilizes a paraphraser adversarially trained on the training dataset. The formats for each attack prompt are described in Figures 9 to 11. Based on the experimental results in Table 4, we confirm that our proposed model demonstrates a non-trivial performance improvement compared to other baselines against these attacks.
<EXAMPLES>: [news article] Determine whether the given <EXAMPLES> content is fact or fake, mentioning the style of the article that led to that decision following format: Answer: {Fact or Fake} Reason: {one-sentence reason }
Your current article is: [news article] But this article is detected as [Answer] due to the following [Reason]. Based on the above information, rewrite a new improved article not to be detected as [news label], maintaining the original content, as follows:
Figure 9: The format of Adversarial prompt. First, use the prompt above to extract the answer and reason of factuality of the given news article. Then, use the following prompt for paraphrasing to avoid detection based on the identified reason. The "news article" section contains the original news text, and the "news label" section contains the corresponding news label.
Summarize the following article, ensuring the content remains the same: [news article]
Figure 10: The format for Summarization prompt. The “news article" part will be filled with the original news text.
<EXAMPLES>: [news article example] Rewrite the following article as the writing style of <EXAMPLES> : [news article]
Figure 11: The format for In-Context prompt. The "news article example" section is filled with an example from a specific publisher, while the "news article" section contains the original news text.
Comparison with LLM-based Baselines.
We compared AdStyle with other LLM-based baselines: GPT-3.5 OpenAI (2022), GPT-4 OpenAI (2023), and In-Context Learning Brown et al. (2020). OpenAI APIs are utilized to predict news article labels, with the temperature set to 0 to ensure stable and consistent predictions. In the case of In-context Learning, one example each of fake news and real news was provided. The prompt formats for the LLM-based baselines are as follows:
Does the following contain real or fake news? Answer in one word with either ‘Real’or ‘Fake’: [news article]
Figure 12: Prompt for LLM-based baselines. The "news article" section contains the original news text.
Table 5: Performance comparison of different LLM-based baselines on the PolitiFact Dataset. (NY: The New York Times, TS: The Sun, NE: National Enquirer)
Model CNN NY TS NE Clean
GPT-3.5 0.5820 0.6242 0.6173 0.5274 0.7037
GPT-4 0.7133 0.6748 0.7025 0.6021 0.7823
In-Context 0.6954 0.6504 0.6875 0.5754 0.7383
Ours 0.9174 0.9444 0.9520 0.9460 0.9460
Full result on style-conversion attack.
Table 6 reports the AUC for each style-conversion attack. Paired sample t-tests were conducted to compare our method with the second-best baseline, Sheepdog, across different ratios for each dataset. Except for the 25% ratio in the Gossipcop dataset, all p-values were significantly smaller than 0.05, indicating statistically significant differences.
Table 6: Performance comparison with AdStyle in four different style—conversion attacks. For the attack scenario, we report the average AUC of four style-conversion attacks. The best results are marked in bold. Our model demonstrates a significant performance improvement over all text-based fake news detectors in diverse style-conversion attack scenarios. (NY: The New York Times, TS: The Sun, NE: National Enquirer )
CNN Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.5664 0.6566 0.8127 0.6841 0.7913 0.8518 0.8563 0.8175 0.8056
UDA 0.5906 0.7174 0.8009 0.7294 0.7859 0.8599 0.8305 0.8594 0.9021
RADAR 0.7205 0.7301 0.7926 0.7683 0.8138 0.8666 0.8099 0.8400 0.8873
ENDEF 0.5426 0.6679 0.7588 0.7288 0.7839 0.8624 0.7695 0.8666 0.8353
SheepDog 0.6305 0.8269 0.8683 0.7472 0.8393 0.8706 0.9200 0.9022 0.9544
AdStyle 0.7547 0.8822 0.9174 0.8195 0.8429 0.8744 0.9344 0.9430 0.9646
NY Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.5899 0.7334 0.8687 0.6750 0.7797 0.8465 0.8522 0.8939 0.8868
UDA 0.6189 0.7644 0.8744 0.7207 0.7795 0.8558 0.8745 0.9184 0.9297
RADAR 0.6966 0.7853 0.8586 0.7481 0.7960 0.8574 0.8701 0.8974 0.9176
ENDEF 0.6152 0.7630 0.8150 0.7227 0.7837 0.8609 0.8341 0.9067 0.8918
SheepDog 0.6310 0.7990 0.9090 0.7420 0.8309 0.8650 0.8937 0.9254 0.9645
AdStyle 0.7599 0.8940 0.9444 0.8118 0.8389 0.8686 0.9243 0.9594 0.9720
TS Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6913 0.6859 0.8789 0.7068 0.7918 0.8430 0.7822 0.8727 0.8974
UDA 0.6646 0.8072 0.8883 0.7488 0.7904 0.8613 0.8662 0.9145 0.9446
RADAR 0.7271 0.7173 0.8909 0.7594 0.8062 0.8620 0.8564 0.8882 0.9132
ENDEF 0.7001 0.8132 0.8444 0.7603 0.7909 0.8639 0.8445 0.8955 0.9043
SheepDog 0.7070 0.8408 0.9096 0.7580 0.8395 0.8688 0.8962 0.9217 0.9665
AdStyle 0.8444 0.9002 0.9520 0.8117 0.8347 0.8751 0.9219 0.9521 0.9727
NE Politifact Gossipcop Constraint
0.1 0.25 1 0.1 0.25 1 0.1 0.25 1
Vanilla 0.6432 0.6856 0.8591 0.7020 0.7875 0.8399 0.7833 0.8853 0.9067
UDA 0.6557 0.7892 0.8622 0.7450 0.7902 0.8595 0.8747 0.9191 0.9423
RADAR 0.7428 0.7269 0.8864 0.7573 0.8026 0.8556 0.8775 0.8919 0.9161
ENDEF 0.6924 0.7877 0.8355 0.7502 0.7897 0.8589 0.8456 0.9112 0.9025
SheepDog 0.6636 0.8268 0.9166 0.7521 0.8331 0.8632 0.8606 0.9257 0.9667
AdStyle 0.8029 0.8913 0.9460 0.8166 0.8391 0.8702 0.9092 0.9579 0.9771